# Generalist World Model Pre-Training for Efficient Reinforcement Learning

### v1

@@ 4.2 WPT Enables Sample-efficient Fine-tuning
Results   Figure 4 shows comparison results with baselines. Our method outperforms or matches all compared baselines by a large margin. Compared to R3M, WPT stresses the importance of world model pre-training as well as reusing offline data during fine-tuning compared to representation learning only. R3M fails to improve sample efficiency on most of the tested tasks. The same conclusion is obtained in Hansen et al. (2023). UDS and ExPLORe reuse offline data by labeling offline data with zero rewards and UCB rewards respectively. The labeled offline data is concatenated with online data for off-policy updates. We find UDS only shows slightly better performance on the Walker Run task compared to R3M and JSRL-BC, showing the ineffectiveness of directly labeling offline data with zero rewards. ExPLORe shows better performance on 2/3 locomotion tasks compared to other baselines and shows signs of meaningful progress on two challenging manipulation tasks. However, WPT still outperforms ExPLORe by a large margin. Compared to ExPLORe, WPT demonstrates the superiority of world model pre-training as well as execution guidance.
Compared to JSRL-BC, our method also demonstrates a clear performance boost. JSRL-BC’s performance heavily depends on the distribution of offline datasets. It can achieve good performance if a good prior actor can be extracted from the offline data. However, in our case, since we aim to lift the assumption of expert trajectories, the pre-trained BC agent usually fails to work well in the target task. As a result, JSRL-BC is only slightly better than the rest baselines on the Cheetah Run Hard, Assembly, and Shelf Place tasks while achieving similar results on the rest. In contrast, WPT works well on leveraging non-expert offline data. For example, the Quadruped Walk shows that WPT can nicely benefit from exploratory offline data collected by unsupervised RL, which enables it to control the Ant robot to walk forward from pixel inputs with only 100 trials.
In Fig. 1 and App. D, we further compare our method with two widely used learning-from-scratch baselines, Dreamerv3 and DrQv2, on 72 tasks. We show that by leveraging the non-curated offline data, WPT can clearly improve the sample efficiency on a wide range of tasks. Furthermore, our method achieves promising performance on several hard exploration tasks, while learning-from-scratch baselines fail.
4.3 WPT Enables Fast Task Adaptation
In many real-world applications, we want the robot to continuously adapt to new tasks. In this section, we investigate whether WPT benefits task adaptation. In this setting, the agent is required to incrementally solve a sequence of tasks. The setting is very similar to continual reinforcement learning (CRL) or life-long RL (Parisi et al., 2019; Khetarpal et al., 2022) but with a limited set of tasks for simplicity. Note that CRL has a broad scope; assumptions and experiment setups vary among methods, which makes it difficult to set up a fair comparison with other methods. Instead of proposing a SOTA CRL method, we aim to demonstrate that our method offers a simple yet general recipe to leverage previous data that also fits the CRL setting.
Setup & baselines   We set our continual multi-task adaption experiment based on the Ant robot from the DeepMind Control Suite. Specifically, the agent needs to sequentially solve stand, walk, run, jump, roll, and roll fast tasks. We train 300K environment steps for each task. To solve one task, the agent accesses all previous experiences as well as the model’s weights trained on previous tasks. To have a fair comparison, i.e., having comparable model parameters and eliminating the potential effects from pre-training on other tasks, we pre-train a small world model only on the Ant domain. We compare our method with PackNet (Mallya & Lazebnik, 2018), which is a common baseline for continual learning. PackNet iteratively prunes the actor’s parameters by keeping the weights with a larger magnitude while re-initializing the rest. By doing so, the actor maintains skills learned previously, which may help the learning of new tasks. For each new task, we fine-tune the actor model by iterative pruning while randomly initializing the critic model as rewards are not shared among tasks.
Results   Figure 5 shows policy performance on incremental task adaption. As we can see WPT outperforms baselines by a large margin. Compared to PackNet, WPT shows great potential for pre-training of world models with non-curated offline data together with experience rehearsal and actor prior. For the tested Ant domain, WPT enables the agent to adapt to new tasks within only 100 trials on all six tasks. With the limited sample budget, PackNet only achieves $\sim 20-60\%$ episodic returns of WPT. We would argue the diversity of the non-curated offline data contributes to WPT’s supreme performance on task adaptation.
4.4 Ablations
In this section, we investigate the role of each component of our method. We use the same set of tasks used in Fig. 4. As shown in Fig. 6, world model pre-training shows promising results when the offline dataset consists of diverse trajectories, such as data collected by exploratory agents, while fails to work well when the offline data distribution is relatively narrow as in the Meta-World tasks. This is due to i) world model pre-training alone failing to fully leverage the rich action prior in the offline data and ii) the distributional shift between offline and online data hurts the world model fine-tuning. The proposed retrieval-based experience rehearsal and execution guidance helps the agent benefit from action prior in the offline data and enriches the online sample diversity, which together enables WPT to achieve strong performance on a wide range of tasks.
5 Conclusion
We propose WPT, a simple yet efficient approach to leverage non-curated offline data. We show that the generalist world model pre-trained on non-curated data can boost RL training spanning multiple tasks and multiple embodiments. Together with retrieval-based experience rehearsal and execution guidance, WPT outperforms baselines on a wide range of tasks, including 22 locomotion tasks and 50 manipulation tasks. WPT unlocks ample sources of offline data and our results show that leveraging this non-curated data leads to strong performance. Nevertheless, our method can be improved in multiple ways, for example, by extending our methods to real-world applications as well as proposing novel world model architectures.
Impact Statement
@@ 5 Conclusion
Appendices
Appendix A Implementation Details
A.1 Behavior Cloning
The Behavior Cloning methods used in both the execution guidance of WPT and JSRL-BC are the same. We use a four-layer convolutional neural network (LeCun et al., 1995) with kernel depth [32, 64, 128, 256] following a three-layer MLPs with LayerNorm (Ba et al., 2016) after all linear layers.
We list the used encoder and actor architectures for reference.
@@ A.1 Behavior Cloning
A.2 JSRL+BC
Jump-start RL (Uchendu et al., 2023) is proposed as an offline-to-online RL method. It includes two policies, a prior policy $\pi_{\theta_{1}}(a|s)$ and a behavior policy $\pi_{\theta_{2}}(a|s)$, where the prior policy is trained via offline RL methods and the behavior policy is updated during the online learning stage. However, offline RL requires the offline dataset to include rewards for the target task. To extract behavior policy from the offline dataset, we use the BC agent described above as the prior policy. During online training, in each episode, we randomly sample the rollout horizon $h$ of the prior policy from a pre-defined array np.arange(0, 101, 10). We then execute the prior policy for $h$ steps and switch to the behavior policy until the end of an episode.
A.3 ExPLORe
@@
Appendix B More Ablation Study
We include per-task results used in Sec. 4.4 in Fig. 7.
Figure 7: Ablation study of the role of each component.
We further compare results using uncertainty-based rewards as suggested in (Li et al., 2023) compared with our expectation guidance. Instead of using UCB for reward labels, we use Optimistic Thompson Sampling (OTS) (Hu et al., 2023b), which is less sensitive w.r.t the selection of hyperparameters. As shown in Fig. 8, our method outperforms the variant using OTS on Assembly and Stick Pull by a large margin, showing the effectiveness of using execution guidance.
Figure 8: Ablation study of using uncertainty-based reward labeling.
Appendix C Algorithm
@@
Algorithm 1 Generalist World Model Pre-Training for Efficient RL
0:  Non-curated offline data $\mathcal{D}_{\text{off}}$, Online data $\mathcal{D}_{\text{on}}\leftarrow\emptyset$, Retrieval data $\mathcal{D}_{\text{retrieval}}\leftarrow\emptyset$              World model $f_{\theta},q_{\theta},p_{\theta},d_{\theta}$              Policy $\pi_{\phi_{\text{RL}}}$, $\pi_{\phi_{\text{BC}}}$, Value function $v_{\phi}$ and Reward $r_{\xi}$.
@@
        Generate imaginary trajectories $\tilde{\tau}=\{s_{t},a_{t},s_{t+1}\}_{t=0}^{T}$ by rolling out $h_{\theta},p_{\theta}$ with policy $\pi_{\phi_{\text{RL}}}$.
@@
Appendix D Full Results
In Table 2 and Table 3, we list the success rate of 50 Meta-World benchmark tasks with pixel inputs. In Table 4, we list the episodic return of DMControl of 22 tasks. We compare WPT at 150k samples with two widely used baselines Dreamer v3 and DrQ v2 at both 150k samples and 1M samples. We mark the best result with a bold front at 150k samples and use underlining to mark the highest score overall.
D.1 Meta-World Benchmark
@@
Tasks Dreamer v3 @ 1M DrQ v2 @ 1M Dreamer v3 @ 150k DrQ v2 @ 150k WPT (ours) @ 150k
@@
More results in the next page, see Table 3
@@
Tasks Dreamer v3 @ 1M DrQ v2 @ 1M Dreamer v3 @ 150k DrQ v2 @ 150k WPT (ours) @ 150k

### v2

@@ Appendix A More Results
Appendix B Theoretical Analysis
In this section, we give a theoretical analysis of the main conclusions in our paper.
B.1 Proof of the Benefits of Experience Retrieval
Proposition 1.
Experience retrieval reduces distribution shift during online fine-tuning, compared to using the full offline dataset directly, in the sense that
$\mathbb{E}_{s\sim p_{\text{retrieved}},\ s_{\text{on}}\sim p_{\text{on}}}\left[||s-s_{\text{on}}||_{2}\right]<\mathbb{E}_{s\sim p_{\text{off}},\ s_{\text{on}}\sim p_{\text{on}}}\left[||s-s_{\text{on}}||_{2}\right].$ (5)
Proof.
Let $p_{\text{off}}(s)$, $p_{\text{on}}(s)$, and $p_{\text{retrieved}}(s)$ denote the state distributions of the non-curated offline dataset $\mathcal{D}_{\text{off}}$, the online dataset $\mathcal{D}_{\text{on}}$, and the retrieved dataset $\mathcal{D}_{\text{retrieved}}\subset\mathcal{D}_{\text{off}}$, respectively. We simplify the notation as
$\mathbb{E}_{s\sim p,\ s_{\text{on}}\sim p_{\text{on}}}\left[||s-s_{\text{on}}||_{2}\right]\quad\text{as}\quad\mathbb{E}_{s\sim p}\left[d(s,s_{\text{on}})\right].$
Since $\mathcal{D}_{\text{retrieved}}\subset\mathcal{D}_{\text{off}}$, the distribution $p_{\text{off}}(s)$ can be expressed as a mixture distribution:
$p_{\text{off}}(s)=\alpha\cdot p_{\text{retrieved}}(s)+(1-\alpha)\cdot p_{\text{rest}}(s),$
where $p_{\text{rest}}(s)$ is the distribution over the remaining offline data, and $\alpha=\frac{|\mathcal{D}_{\text{retrieved}}|}{|\mathcal{D}_{\text{off}}|}$ denotes the fraction of samples in the retrieved dataset.
The expected total variation for the mixture distribution decomposes as:
$\mathbb{E}_{s\sim p_{\text{off}}}\left[d(s,s_{\text{on}})\right]=\alpha\cdot\mathbb{E}_{s\sim p_{\text{retrieved}}}\left[d(s,s_{\text{on}})\right]+(1-\alpha)\cdot\mathbb{E}_{s\sim p_{\text{rest}}}\left[d(s,s_{\text{on}})\right].$ (6)
Assume that $\mathcal{D}_{\text{retrieved}}$ is constructed by selecting states such that $\|s_{\text{retrieved}}-s_{\text{on}}\|<\epsilon$, for some small $\epsilon>0$. Consequently, states in $\mathcal{D}_{\text{rest}}$ satisfy $\|s_{\text{rest}}-s_{\text{on}}\|\geq\epsilon$. This construction implies the following bounds:
$\mathbb{E}_{s\sim p_{\text{retrieved}}}\left[d(s,s_{\text{on}})\right]$ $<\epsilon^{\prime},$ (7)
$\mathbb{E}_{s\sim p_{\text{rest}}}\left[d(s,s_{\text{on}})\right]$ $\geq\epsilon^{\prime},$ (8)
for some $\epsilon^{\prime}>0$. Therefore, it follows that
$\mathbb{E}_{s\sim p_{\text{rest}}}\left[d(s,s_{\text{on}})\right]>\mathbb{E}_{s\sim p_{\text{retrieved}}}\left[d(s,s_{\text{on}})\right].$
Substituting into Equation (6) yields:
$\mathbb{E}_{s\sim p_{\text{off}}}\left[d(s,s_{\text{on}})\right]$ $=\alpha\cdot\mathbb{E}_{s\sim p_{\text{retrieved}}}\left[d(s,s_{\text{on}})\right]+(1-\alpha)\cdot\mathbb{E}_{s\sim p_{\text{rest}}}\left[d(s,s_{\text{on}})\right]$
$>\alpha\cdot\mathbb{E}_{s\sim p_{\text{retrieved}}}\left[d(s,s_{\text{on}})\right]+(1-\alpha)\cdot\mathbb{E}_{s\sim p_{\text{retrieved}}}\left[d(s,s_{\text{on}})\right]$
$=\mathbb{E}_{s\sim p_{\text{retrieved}}}\left[d(s,s_{\text{on}})\right].$
Thus, the expected total variation between the retrieved data and online data is strictly smaller than that between the full offline data and online data. ∎
Proposition 2.
Experience retrieval helps prevent catastrophic forgetting during online fine-tuning.
Definition 1 (Catastrophic Forgetting due to Data Distribution Shift).
Catastrophic forgetting occurs when a neural network, after training on a new data distribution, experiences a significant performance drop on previously learned tasks due to the overwriting of representations from earlier distributions, caused by biased parameter updates towards the new distribution.
Proof.
Following the previous notations, let $\mathcal{D}_{\text{on}}$ and $\mathcal{D}_{\text{retrieved}}$ denote the online dataset and the retrieved offline dataset, respectively. The objective in Eq. 1 can be written as:
$\mathcal{L}_{\text{mixed}}(\theta)$ $=\mathcal{L}_{\text{on}}(\theta)+\lambda\cdot\mathcal{L}_{\text{retrieved}}(\theta)$
$=\mathbb{E}_{p_{\theta},q_{\theta},\ (o,a)\sim\mathcal{D}_{\text{on}}}\left[\sum_{t=1}^{T}-\ln p_{\theta}(o_{t}\,|\,z_{t},h_{t})+\beta\cdot\mathrm{KL}\left(q_{\theta}(z_{t}\,|\,h_{t},o_{t})\ \|\ p_{\theta}(z_{t}\,|\,h_{t})\right)\right]$
$+\lambda\cdot\mathbb{E}_{p_{\theta},q_{\theta},\ (o,a)\sim\mathcal{D}_{\text{retrieved}}}\left[\sum_{t=1}^{T}-\ln p_{\theta}(o_{t}\,|\,z_{t},h_{t})+\beta\cdot\mathrm{KL}\left(q_{\theta}(z_{t}\,|\,h_{t},o_{t})\ \|\ p_{\theta}(z_{t}\,|\,h_{t})\right)\right].$
Assuming the $\lambda$ is a monotonic function of $\alpha=\frac{|\mathcal{D}_{\text{retrieved}}|}{|\mathcal{D}_{\text{off}}|}$ and $\lambda>0$, since $\mathcal{D}_{\text{retrieved}}\subset\mathcal{D}_{\text{off}}$, the term $\mathcal{L}_{\text{retrieved}}(\theta)$ acts as a regularizer during online updates, constraining parameter changes on $\mathcal{D}_{\text{on}}$ in a way that preserves performance on the retrieved offline distribution $p_{\text{retrieved}}$. This mitigates the risk of catastrophic forgetting by anchoring the model to previously seen data. ∎
B.2 Proof of Improved Performance with Execution Guidance
Proposition 3 (Performance Improvement via Execution Guidance).
Let $\pi^{e}$ denote an exploration policy and $\pi^{g}$ a guide policy obtained via imitation learning. Assume that $\pi^{g}$ outperforms $\pi^{e}$ at the early stage of training, that is, there exists a constant $\delta>0$ such that for all states $s$ in the support of the state visitation distribution:
$\mathbb{E}_{a\sim\pi^{g}(\cdot|s)}[A_{\pi^{e}}(s,a)]\geq\delta$ (9)
Let $\tilde{\pi}$ be an $\alpha$-coupled policy (execution guidance) derived from $\pi^{e}$ and $\pi^{g}$, as defined in Definition 1 of [68], such that:
$P(a\neq a_{g}|s)\leq\alpha,$ (10)
where $a\sim\pi^{e}(\cdot|s)$ and $a_{g}\sim\pi^{g}(\cdot|s)$.
Then, the performance of the $\alpha$-coupled policy $\tilde{\pi}$ exceeds that of the exploration policy $\pi^{e}$ by at least:
$\eta(\tilde{\pi})-\eta(\pi^{e})\geq\frac{\alpha\cdot\delta}{1-\gamma},$ (11)
where $\gamma\in[0,1)$ is the discount factor.
Proof.
By Lemma 1 in Trust Region Policy Optimization (TRPO) [68], the difference in policy performance can be expressed as:
$\eta(\tilde{\pi})=\eta(\pi^{e})+\mathbb{E}_{\tau\sim\tilde{\pi}}\left[\sum_{t=0}^{\infty}\gamma^{t}A_{\pi^{e}}(s_{t},a_{t})\right]$ (12)
Define the expected advantage at state $s$ as:
$\bar{A}_{\pi^{e}}(s)=\mathbb{E}_{a\sim\tilde{\pi}(\cdot|s)}[A_{\pi^{e}}(s,a)]$
Given the definition of an $\alpha$-coupled policy, we can decompose this expectation:
$\bar{A}_{\pi^{e}}(s)$ $=(1-\alpha)\cdot\mathbb{E}_{a\sim\pi^{e}(\cdot|s)}[A_{\pi^{e}}(s,a)]+\alpha\cdot\mathbb{E}_{a\sim\pi^{g}(\cdot|s)}[A_{\pi^{e}}(s,a)]$ (13)
Since $\mathbb{E}_{a\sim\pi^{e}(\cdot|s)}[A_{\pi^{e}}(s,a)]=0$ (by definition of the advantage function), we have:
$\bar{A}_{\pi^{e}}(s)$ $=\alpha\cdot\mathbb{E}_{a\sim\pi^{g}(\cdot|s)}[A_{\pi^{e}}(s,a)]$
$\geq\alpha\cdot\delta$ (14)
The policy performance difference can then be bounded as:
$\eta(\tilde{\pi})-\eta(\pi^{e})$ $=\mathbb{E}_{\tau\sim\tilde{\pi}}\left[\sum_{t=0}^{\infty}\gamma^{t}\bar{A}_{\pi^{e}}(s_{t})\right]$
$=\sum_{s}\rho_{\tilde{\pi}}(s)\bar{A}_{\pi^{e}}(s)$
$\geq\sum_{s}\rho_{\tilde{\pi}}(s)\cdot\alpha\cdot\delta$
$=\alpha\cdot\delta\cdot\sum_{s}\rho_{\tilde{\pi}}(s),$ (15)
where $\rho_{\tilde{\pi}}(s)$ is the discounted state visitation frequency under $\tilde{\pi}$.
Since $\sum_{s}\rho_{\tilde{\pi}}(s)=\frac{1}{1-\gamma}$ (the sum of discounted state visitation frequencies equals the effective horizon length), we have:
$\eta(\tilde{\pi})-\eta(\pi^{e})$ $\geq\frac{\alpha\cdot\delta}{1-\gamma}$ (16)
This establishes that the performance improvement of the $\alpha$-coupled policy $\tilde{\pi}$ over the exploration policy $\pi^{e}$ is at least $\frac{\alpha\cdot\delta}{1-\gamma}$, which is strictly positive when $\alpha>0$, $\delta>0$, and $\gamma<1$. ∎
Appendix C More Related Work
Unsupervised RL   In unsupervised RL, an agent explores the environment based on intrinsic motivations, and the models’ parameters are initialized during this self-motivated exploration stage, aiming for fast downstream task learning [28, 62, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78]. Our problem setting differs from unsupervised RL in several ways: i) Unsupervised RL interacts with the environment actively while we leverage static offline datasets, ii) unsupervised RL gives a specific focus on designing different intrinsic rewards, while our setting focuses on improving sample efficiency by leveraging unlabeled datasets.
Generalist Agents   RL methods usually perform well on a single task [79, 80], however, this contrasts with humans who can perform multiple tasks well. Recent works have proposed generalist agents that master a diverse set of tasks with a single agent [81, 82, 83, 84]. These methods typically resort to scalable models and large datasets and are trained via imitation learning [82, 85, 86, 87]. In contrast, we train a generalist world model and use it to boost RL performance for multiple tasks and embodiments.
World models   World models learn to predict future observations or states based on historical information. World models have been widely investigated in online model-based RL [54, 88, 89, 90, 91]. Recently, the community has started investigating scaling world models [88], for example, Hu et al. [92], Pearce et al. [93], Wu et al. [26], Agarwal et al. [94] train world models with Diffusion Models or Transformers. However, these models are usually trained on demonstration data. In contrast, we explore the offline-to-online RL setting – closely fitting the pre-train and then fine-tune paradigm – and we focus on leveraging reward-free and multi-embodiment data to increase the amount of available data for pre-training. We further identify the distributional shift issue when fine-tuning the pre-trained world model and mitigate the issue by proposing experience rehearsal and execution guidance.
Appendix D Limitations
Although demonstrating strong performance on a diverse set of tasks, our method has the following limitations. 1) The world model architecture used in our paper is the recurrent state space model. This model is built upon RNN, which can be limited for scaling. This can be mitigated by using a Transformer and a diffusion-based world model. However, we note that the main conclusion of this paper should still be valid. 2) We do not thoroughly discuss the generalization ability of the pre-trained world model. With DMControl tasks, our method shows a promising trend in generalizing to unseen tasks. However, generalization to new embodiments or novel configurations is still challenging, which requires even diverse training data. 3) The non-curated offline data used in our paper, although lifting several key assumptions in previous offline-to-online RL, is still in-domain data, i.e., our current method is not able to leverage the vast in-the-wild data. A promising direction is to combine in-the-wild data for pre-training as in [26] and the domain-specific “in-house” data (as used in our paper) for post-training. 4) We only conduct experiments in the simulator. Considering the sample efficiency of our proposed method, it could be promising to conduct experiments on real-world applications.
Appendix E Impact Statement
@@
Appendix F Compute Resources
We conduct all experiments on clusters equipped with AMD MI250X GPUs, 64-core AMD EPYC “Trento" CPUs, and 64 GBs DDR4 memory. For pre-training, it takes $\sim 48$ GPU hours for 150k steps. For fine-tuning, it tasks $\sim 8$ GPU hours per run for 150K environment steps. Note that due to AMD GPUs not supporting hardware rendering, the training time should be longer than using Nvidia GPUs. To reproduce the GSA’s results in Fig. 3, it roughly takes 8 h * 72 tasks * 3 seeds = 1728 GPU hours.
Appendix G Implementation Details
G.1 Behavior Cloning
The Behavior Cloning methods used in both the execution guidance of GSA and JSRL-BC are the same. We use a four-layer convolutional neural network [95] with kernel depth [32, 64, 128, 256] following a three-layer MLPs with LayerNorm [96] after all linear layers.
We list the adopted encoder and actor architectures for reference.
@@ G.1 Behavior Cloning
G.2 JSRL+BC
Jump-start RL [33] is proposed as an offline-to-online RL method. It includes two policies, a prior policy $\pi_{\theta_{1}}(a|s)$ and a behavior policy $\pi_{\theta_{2}}(a|s)$, where the prior policy is trained via offline RL methods and the behavior policy is updated during the online learning stage. However, offline RL requires the offline dataset to include rewards for the target task. To extract behavior policy from the offline dataset, we use the BC agent described above as the prior policy. During online training, in each episode, we randomly sample the rollout horizon $h$ of the prior policy from a pre-defined array np.arange(0, 101, 10). We then execute the prior policy for $h$ steps and switch to the behavior policy until the end of an episode.
G.3 ExPLORe
@@
Appendix H Algorithm
@@
Algorithm 1 Efficient RL by Guiding Generalist World Models with Non-Curated Offline Data
Non-curated offline data $\mathcal{D}_{\text{off}}$, Online data $\mathcal{D}_{\text{on}}\leftarrow\emptyset$, Retrieval data $\mathcal{D}_{\text{retrieval}}\leftarrow\emptyset$      World model $f_{\theta},q_{\theta},p_{\theta},d_{\theta}$      Policy $\pi_{\phi_{\text{RL}}}$, $\pi_{\phi_{\text{BC}}}$, Value function $v_{\phi}$ and Reward $r_{\xi}$.
@@
         Generate imaginary trajectories $\tilde{\tau}=\{s_{t},a_{t},s_{t+1}\}_{t=0}^{T}$ by rolling out $h_{\theta},p_{\theta}$ with $\pi_{\phi_{\text{RL}}}$.
@@
Appendix I Full Results
In Table 2 and Table 3, we list the success rate of 50 Meta-World benchmark tasks with pixel inputs. In Table 4, we list the episodic return of DMControl of 22 tasks. We compare GSA at 150k samples with two widely used baselines Dreamer v3 and DrQ v2 at both 150k samples and 1M samples. We mark the best result with a bold front at 150k samples and use underlining to mark the highest score overall.
I.1 Meta-World Benchmark
@@
Tasks Dreamer v3 @ 1M DrQ v2 @ 1M Dreamer v3 @ 150k DrQ v2 @ 150k GSA @ 150k
@@
More results see Table 3
@@
Tasks Dreamer v3 @ 1M DrQ v2 @ 1M Dreamer v3 @ 150k DrQ v2 @ 150k GSA (ours) @ 150k